OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (anonymous):

are u looking for the slope? or the y-mx+b?

OpenStudy (anonymous):

The x coordinate.

OpenStudy (anonymous):

what the answer they give u on it

OpenStudy (anonymous):

They didnt give me any

OpenStudy (anonymous):

oh ok will I got slope is -1

OpenStudy (anonymous):

I dont get it. So whats the x coordinate??

OpenStudy (anonymous):

@jim_thompson5910 can you help??

jimthompson5910 (jim_thompson5910):

first you need to find the equation of the line through the two points

jimthompson5910 (jim_thompson5910):

are you able to do that?

OpenStudy (anonymous):

nope. can you bring me through the process?

jimthompson5910 (jim_thompson5910):

do you know how to find the slope of the line through those two points?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

have you seen the slope formula m = (y2-y1)/(x2-x1) before?

OpenStudy (anonymous):

ohhh yes I have. Ill try to solve for it just a sec

jimthompson5910 (jim_thompson5910):

ok plug in the given points and tell me what slope you get

OpenStudy (anonymous):

6/6 = 1

OpenStudy (anonymous):

so @jim_thompson5910 the slope is one tho

jimthompson5910 (jim_thompson5910):

not quite m = (y2-y1)/(x2-x1) m = ( -3- 3)/(1-(-5)) m = ( -3- 3)/(1+5) m = (-6)/6 m = -1

jimthompson5910 (jim_thompson5910):

You lost a sign somewhere

OpenStudy (anonymous):

oh yup I did :/ so now whats next??

jimthompson5910 (jim_thompson5910):

now use this to find the y-intercept

jimthompson5910 (jim_thompson5910):

y = mx+b y = -1*x + b 3 = -1*(-5) + b 3 = 5 + b b = ??

OpenStudy (anonymous):

b=-2

jimthompson5910 (jim_thompson5910):

so the equation is y = -x - 2

jimthompson5910 (jim_thompson5910):

to find the x-intercept, you plug in y = 0 and solve for x

OpenStudy (anonymous):

x=-2

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

yayyy:)) thank you can you help me with one more? Mary is 14 years younger than her sister Jane. In four years, Jane will be 3 times older than Mary. How old is Jane now? 3 7 17 21

jimthompson5910 (jim_thompson5910):

how far did you get here

OpenStudy (anonymous):

m+14=j

jimthompson5910 (jim_thompson5910):

Mary is 14 years younger than her sister Jane means that M = J + 14

OpenStudy (anonymous):

that wouldnt really be possible cause jane is older than mary

jimthompson5910 (jim_thompson5910):

oh wait, misread that, it should be M = J - 14 sorry about that

jimthompson5910 (jim_thompson5910):

so if you rearrange terms you get J = M + 14

jimthompson5910 (jim_thompson5910):

in 4 years, Jane is J+4 years old in 4 years, Mary is M+4 years old

OpenStudy (anonymous):

m=j-14 j+4=3(m+4) solving we get mary =3 jane=17

jimthompson5910 (jim_thompson5910):

good, you got it